OpenStudy (darkigloo):

How would I find the area under the curve withouth plotting the points? (calculus) f(x)=sinx x: 0, pi/4, pi/2, 3pi/4, pi f(X): 0, √2/2, 1, √2/2, 0

OpenStudy (anonymous):

integrate it,be carefull with negative values

OpenStudy (darkigloo):

you mean \[\int\limits_{0}^{\pi}sinxdx\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2

OpenStudy (darkigloo):

i got cos(pi) - cos(0) = -2

OpenStudy (johnweldon1993):

Well the integral of sin = -cos -cos(pi) + cos(0) = ? -(-1) + 1 = ? 1 + 1 = 2

OpenStudy (anonymous):

no the integral is -cos(x)

OpenStudy (darkigloo):

right! i just realized that

OpenStudy (darkigloo):

so is that RRAM?

OpenStudy (darkigloo):

or LRAM?

OpenStudy (anonymous):

what is RRAM?

OpenStudy (darkigloo):

lram is left hand, rram is right hand. i know that when you do left hand, you dont include the last y value, and when you do right hand, you don't include the first y value.